• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 08:08:06 on February19,2007

    Entry number 194171

    keyword=Owl shift summary

    Shift crew: R. Feuerbach, J. Miller

    Shift summary



    Conditions: Beam energy = 362.26 MeV/c (readback fluctuating by +/- 0.1 MeV), LHRS at 35deg, RHRS at 24deg.

    0:05 Finished runs 6571/26555. Moving to B4C.
    1:25 Found run 26560 had no events, so runs 6575/26560 were declared 'junk'
    2:40 Restarted Right arm CODA.
    3:10 Moved to kinematic DD at 39deg.
    3:12 MCC says they can only give us up to ~15 microamps, instead of the 20 that we want for Ta.

    3:30 Beam position adjusted to be back close to 0 in X (was out at +.6 at BPMB). Starting new runs with the improved position.

    7:45 Moved LHRS to 43 degrees.





    Run Summary



    Run Kin Target Start Comment
    6571/26555 CC Li 23:48 1 microamp
    6572/26556 CC B4C 0:09 1 microamp
    6573/26557 CC B4C 0:28 1 microamp
    6574/26558 CC C 0:53 1.5 microamp
    6575/26560 CC C 1:15 1.3 microamp -- junk
    6576/26561 CC C 1:25 1.3 microamp
    6577/26562 CC Ta 1:53 10.8 microamp
    6579/26565 CC Ta 2:40 15 microamp
    LHRS at 39deg.
    6580/26566 DD Ta 3:13 16 microamp, adjust beam pos
    6581/26567 DD Ta 3:30 16 microamp
    6582/26568 DD Ta 4:17 16 microamp
    6583/26569 DD C 5:08 3.5 microamp
    6584/26570 DD C 5:31 3.5 microamp
    6585/26571 DD B4C 5:56 1 microamp, DT~15%
    6586/26572 DD B4C 6:08 1.4 microamp/ps3=2
    6587/26573 DD B4C 6:27 1.4 microamp
    6588/26574 DD Li 6:52 1.0 microamp
    6589/26575 DD Li 7:16 2.4 microamp/ps3=2
    LHRS at 43deg.
    6592/26578 EE Li 8:04 2.0 microamp